OpenStudy (lovelyanna):

Refer to the following conditional statement: If a and b are odd integers, then a – b is an even integer. Identify the hypothesis. A. a – b is not an even integer B. a and b are not odd integers C. a and b are odd integers D. a – b is an even integer

OpenStudy (lovelyanna):

I think its C

OpenStudy (lovelyanna):

@whpalmer4

OpenStudy (anonymous):

Yes that is correct

OpenStudy (lovelyanna):

Ok, thanks

OpenStudy (lovelyanna):

Refer to the following conditional statement: If a and b are odd integers, then a • b is an odd integer. Which is the inverse of the statement? A. a • b is an odd integer B. If a • b is an odd integer, then a and b are odd integers. C. If a and b are not odd integers, then a • b is not an odd integer. D. a and b are odd integers I think it's C.

OpenStudy (lovelyanna):

@BlackLabel

OpenStudy (anonymous):

Yup that is correct

OpenStudy (anonymous):

If p then q Inverse: If ~p then ~q

OpenStudy (lovelyanna):

Oh, then B? O.O

OpenStudy (lovelyanna):

Wait, thats not right

OpenStudy (lovelyanna):

Ya, its C

OpenStudy (lovelyanna):

Right?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

You got it right I was just giving you the definition of an inverse statement

OpenStudy (lovelyanna):

Ok, thanks! Can you help with a few more?

OpenStudy (anonymous):

Ya np. Just working at the same time so it may take a lil longer to respond

OpenStudy (lovelyanna):

Ok, thanks! Refer to the following conditional statement: If a and b are odd integers, then a • b is an odd integer. Which is the contrapositive of the statement? A. If a • b is not an odd integer, then a and b are not odd integers. B. If a • b is an odd integer, then a and b are odd integers. C. a • b is an odd integer D. a and b are odd integers

OpenStudy (lovelyanna):

I would say B

OpenStudy (lovelyanna):

Oh wait.. actually, A

OpenStudy (anonymous):

Yup it is A

OpenStudy (lovelyanna):

Sorry! XP

OpenStudy (lovelyanna):

What kind of reasoning is Kim using? Kim notices that 24 = 6 + 18, 63 = 9 + 54, 39 = 12 + 27, and 72 = 9 + 63, so she thinks it is likely that she can write every number divisible by 3 as the sum of two numbers divisible by 3. A. deductive B. inductive C. hypothesis D.conclusion

OpenStudy (lovelyanna):

B?

OpenStudy (anonymous):

Yup I guess we can say inductive

OpenStudy (anonymous):

That is the most apropo even though we usually can not induce like that when proving

OpenStudy (lovelyanna):

Ok, thanks! What kind of reasoning is used here? If a = b and c = d, then a • c = b • d. A. hypothesis B. conclusion C. deductive D. inductive I think this is C

OpenStudy (lovelyanna):

@BlackLabel

OpenStudy (anonymous):

I think so too since this is more general

OpenStudy (lovelyanna):

Yup, thanks! How about this one? Which is a counterexample that disproves the following statement? For all integers x, 1/x \[\le\] x A. x = 5 B. x = 100 C. x = 2 D. x = 0.5

OpenStudy (lovelyanna):

Oh, my bad, the: \[\le\] Should be after 1/x and before x, like this: 1/x <_ x

OpenStudy (lovelyanna):

I'm not sure about this question.. can you explain?

OpenStudy (lovelyanna):

Maybe D..?

OpenStudy (anonymous):

Ok basically the statement is saying that for any integers x then \( \frac{1}{x}\) will always be smaller than or equal to x Now we need to find an example where this is not the case BUt the statement is true for 100 since 1/100<100 and 1/5<5 and 2< 1/2 Now the issue with .5 is that .5 is not an integer. .5 is a fraction and the question specifically states that x must be ab integer

OpenStudy (anonymous):

1/.5=2 so .5<1/.5 But .5 is not an integer So D would be my preferred option eventhough its not really true

OpenStudy (lovelyanna):

Ah, I see! Then D it is?

OpenStudy (anonymous):

It really isnt D but just go for it anyways

OpenStudy (lovelyanna):

Btw, did you realize that the question says: Which is a counterexample that DISPROVES the following statement?

OpenStudy (lovelyanna):

Do, if you say it isn't D, then for this question, it IS D?

OpenStudy (anonymous):

Yaaaa I got that None of them disprove the statement In order to disprove the statement The hypothesis must be true which is that x must be an integer. Since x is not an integer then x can not qualify to disprove the statement

OpenStudy (lovelyanna):

I mean So*

OpenStudy (lovelyanna):

Then its still D?

OpenStudy (anonymous):

Lets go for D. They may have worded the question in correctly and meant to say ration number instead of integer

OpenStudy (anonymous):

rational number*

OpenStudy (lovelyanna):

Alright then! Thanks What is the reason for Statement 2? A. addition property of equality B. property of opposites C. identity property of addition D. associative property of addition

OpenStudy (lovelyanna):

Is it B?

OpenStudy (anonymous):

addition property of opposites The property that states the sum of a number and its opposite is always zero.

OpenStudy (anonymous):

Yup that is what I was thinking

OpenStudy (anonymous):

Only one that makes most sense

OpenStudy (lovelyanna):

Ok, thanks :) This one: Write the contrapositive of the conditional statement and then determine if the conditional statement and the contrapositive are true or false. If 6x ≠ 24, then x ≠ 4. A. If 6x = 24, then x = 4. The conditional statement and the contrapositive are both true. B. If x = 4, then 6x = 24. The conditional statement and the contrapositive are both true. C. If x = 4, then 6x = 24. The conditional statement is false, and the contrapositive is true. D. If x ≠ 4, then 6x ≠ 24. The conditional statement and the contrapositive are both false.

OpenStudy (lovelyanna):

Oh, sorry, its all mixed up O.O A. If 6x = 24, then x = 4. The conditional statement and the contrapositive are both true. B. If x = 4, then 6x = 24. The conditional statement and the contrapositive are both true. C. If x = 4, then 6x = 24. The conditional statement is false, and the contrapositive is true. D. If x ≠ 4, then 6x ≠ 24. The conditional statement and the contrapositive are both false.

OpenStudy (lovelyanna):

I think B

OpenStudy (lovelyanna):

My bad, I meant A

OpenStudy (anonymous):

Im not sure Im flirting btwn A and B Lemme think abt this

OpenStudy (lovelyanna):

Ok

OpenStudy (anonymous):

"A" definitely sounds more correct. So I would go with A but maybe post another question and get someone to take a look at this

OpenStudy (lovelyanna):

Aren't A and B almost the same thing? I mean, all they did was switch these around: If 6x = 24, then x = 4.

OpenStudy (lovelyanna):

Ok, thanks :) I'll do that!

OpenStudy (anonymous):

I would like to disagree the if part is the hypothesis: "A very closely related logical meaning is that for any statement or claim of the form "if A, then B", A is said to the be hypothesis of the claim. This usage transfers the idea of a "hypothesis" from the argument that establishes A⇒B to the naked assertion that A⇒B. "Assumption" is possible here, but appears to be less common than "hypothesis", especially if the claim is written symbolically rather than a theorem statement in prose."

OpenStudy (whpalmer4):

You're right, that's the formal logic usage, the hypothesis being the antecedent. I withdraw my objection, though I'll continue to think that's a ridiculous hypothesis :-)